# Finding Most Shattering Minimum Vertex Cuts of Polylogarithmic Size in Near-Linear Time

### v1

@@ 6.2 High-Degree
Proof of Lemma˜6.9.
Correctness and time complexity of Algorithm˜3 directly follow from Lemmas˜6.10 and 6.11. ∎
@@ 7 Capturing and Listing Unbalanced $k$-Shredders
In the previous sections, we showed an algorithm that takes as input a tuple $(x,\nu,\Pi)$, and lists all $k$-shredders that are captured by $(x,\nu,\Pi)$ as well as an unverified set. We then showed how to resolve unverified sets using casework on the structural properties of $k$-shredders. There is one piece of the puzzle left: the method for capturing unbalanced $k$-shredders. For this, we will leverage random sampling using Lemma˜6.7 as well as recent developments in local flow algorithms as in [FNS+20].
@@ 7.1 Leveraging Local Flow Algorithms
Theorem 4.1 of [FNS+20] only states that their algorithm returns a vertex cut. But they also construct the set of paths $\Pi$. Their algorithm is simple, and we will briefly explain it here. First, we perform the standard vertex-splitting reduction and reduce the problem to finding directed edge-disjoint paths instead. To find the first path, we use DFS starting from a vertex $x$ and explore $k\nu$ volume. Then, we sample a random endpoint $y_{1}$ among all explored edges. Note that $y_{1}$ is in $L$ with probability at least 11/k1-1/k since $\mathrm{vol}(R)\leq\nu$. We reverse the direction of edges on the path from $x$ to $y_{1}$ in the DFS tree and obtain a “residual” graph. Then, we repeat the process in the residual graph to construct the next path from $x$ to $y_{2}$. After $k$ iterations, the endpoints $y_{1},\dots,y_{k}$ of these $k$ paths are in $L$ with probability $(1-1/k)^{k}\geq 1/4$. The paths in the residual graphs can be decomposed via flow decomposition into $k$ directed edge-disjoint paths in the original graph whose total length is $k^{2}\nu$. Finally, these paths correspond to $k$ openly-disjoint paths by the standard reduction in the beginning.
The output of the algorithm described in ˜7.2 directly corresponds to Definition˜2.4. Let $R$ denote the union of all components in $\mathcal{R}$. Notice that $(C,S,R)$ forms a vertex cut such that $\mathrm{vol}(R)<\frac{m}{k}$. The idea is to obtain a seed vertex $x\in R$ using a linear amount of random samples. For each sample, we can directly apply $\texttt{LocalVC}(\cdot,\cdot)$ to obtain the desired set of paths. Furthermore, we can boost the success rate of the algorithm by repeating it a polylogarithmic number of times. In the following section, we formalize this idea.
@@ 7.2 The Algorithm for Unbalanced $k$-Shredders
We first give a high level outline for listing unbalanced $k$-shredders. Let $S$ be an unbalanced $k$-shredder with partition $(C,\mathcal{R})$. With geometric sampling and Lemma˜6.7, we will sample a vertex $x\in\mathcal{R}$ with a volume parameter $\nu$ that satisfies $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. Then, we use ˜7.2 to obtain a set of $k$ openly-disjoint paths $\Pi$, each starting from $x$ and ending at a vertex in $C$ such that the sum of lengths over all paths is at most $k^{2}\nu$. We now have a tuple $(x,\nu,\Pi)$ that captures $S$. After capturing $S$, we call Algorithm˜1 to list $S$ as a $k$-shredder or an unverified set. In the latter case, we can verify whether $S$ is a low-degree $k$-shredder using Algorithm˜2. If this verification step fails, $S$ must be a high-degree $k$-shredder. In this case, we can add $S$ to a global list of unverified sets. This list will be processed after all unbalanced $k$-shredders have been captured. Lastly, we can extract all high-degree $k$-shredders from the list of unverified sets using Algorithm˜3. Pseudocode describing this process is given in Algorithm˜4.
@@
1: $G=(V,E)$ - an undirected $k$-vertex-connected graph
2: $\mathcal{L}$ - a list containing all unbalanced $k$-shredders of $G$
3: $\mathcal{L}\leftarrow\varnothing$
4: $\mathcal{U}\leftarrow\varnothing$
5:
$\Phi\leftarrow$ initialize a pairwise connectivity oracle as in [Kos23]
6:
7: for $i\leftarrow 0$ to $\lceil\log\left(\frac{m}{k}\right)\rceil$ do
8:   $\nu\leftarrow 2^{i}$
9:   for $400\frac{m}{\nu}\log n$ times do
10:    independently sample an edge $(x,y)$ uniformly at random
11:    for $300\log n$ times do
12:      $\Pi\leftarrow\texttt{LocalVC}(x,\nu)$ as in [FNS+20]
13:      $(\mathcal{L}_{local},U)\leftarrow$ call Algorithm˜1 on input $(x,\nu,\Pi)$
14:      $\mathcal{L}\leftarrow\mathcal{L}\cup\mathcal{L}_{local}$
15:      if $U\neq\varnothing$ then
16:       call Algorithm˜2 on input $(x,\nu,\Pi,U)$
17:       if Algorithm˜2 returned true then
18:         $\mathcal{L}\leftarrow\mathcal{L}\cup\{U\}$
19:       else
20:         $\mathcal{U}\leftarrow\mathcal{U}\cup\{(x,\nu,\Pi,U)\}$
21:
22: $\mathcal{L}_{high-degree}\leftarrow$ call Algorithm˜3 on input $\mathcal{U}$
23: $\mathcal{L}\leftarrow\mathcal{L}\cup\mathcal{L}_{high-degree}$
24: return $\mathcal{L}$
@@
Let $S$ be an unbalanced $k$-shredder. Then $S$ is captured by a tuple $(x,\nu,\Pi)$ after line 12 during some iteration of the loop on line 7 with probability $1-n^{-99}$.
@@
Suppose $S$ has partition $(C,\mathcal{R})$. Consider the iteration of the loop on line 7 where $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. By Lemma˜6.7, we will sample an edge $(x,y)$ such that $x\in\mathcal{R}$ with probability $1-n^{-100}$ during the for loop on line 9. The only missing piece of the tuple is the set of paths. This piece is effectively solved by ˜7.2. Let $R$ be the union of all components in $\mathcal{R}$. Notice that $(C,S,R)$ is a vertex cut such that $\mathrm{vol}(R)<\frac{m}{k}$. We can call $\texttt{LocalVC}(x,\nu)$ to obtain a set $\Pi$ of $k$ openly-disjoint paths that start from $x$ and end in $C$ such that the sum of lengths over all paths is at most $k^{2}\nu$ with probability $1/4$. This probability is boosted by the for loop on line 11. So the probability that we fail to obtain such a set of paths over all trials is reduced to $(3/4)^{300\log n}\leq(1/2)^{100\log n}=n^{-100}$. Hence, the joint probability that we sample a vertex $x\in\mathcal{R}$ and obtain the desired set of paths $\Pi$ is at least $(1-n^{-100})^{2}\geq 1-2n^{-100}\geq 1-n^{-99}$ for all $n\geq 2$. ∎
@@
Let $S$ be a $k$-shredder captured by the tuple $(x,\nu,\Pi)$ after line 12. If $S$ is low-degree, then $S\in\mathcal{L}$ after line 18. Otherwise, if $S$ is high-degree, then $S\in\mathcal{L}$ after line 23 with probability $1-n^{-100}$.
@@
Suppose $S$ has partition $(C,\mathcal{R})$ and is captured by $(x,\nu,\Pi)$. If Algorithm˜1 returns $S$ as a $k$-shredder, we are immediately done by lines 13-14. So let us assume that $S$ is reported as an unverified set. We have that $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. If $S$ is low-degree, we have that there exists a vertex $s\in S$ such that $\deg(s)\leq\nu$. By Lemma˜6.3, Algorithm˜2 will return true on input $(x,\nu,\Pi,S)$. Then, $S$ is added to $\mathcal{L}$ by line 18. Otherwise, suppose that $S$ is high-degree. Then, $(x,\nu,\Pi,S)$ is added to the set $\mathcal{U}$ by line 20. By Lemma˜6.9, Algorithm˜3 will return a list of high-degree $k$-shredders containing $S$ on line 22 with probability $1-n^{-100}$. Hence, $S\in\mathcal{L}$ by line 23 with probability $1-n^{-100}$. ∎
@@
There are only three lines of the algorithm where we add sets to $\mathcal{L}$: 13, 18, 23. Line 13 only adds $k$-shredders due to Lemma˜2.5. Line 18 only adds $k$-shredders due to Lemma˜6.3. Lastly, line 23 only adds $k$-shredders due to Lemma˜6.9. ∎
@@
Algorithm˜4 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ time.
@@
Firstly, line 5 runs in $\mathcal{O}(km\log n)$ time by ˜6.2. We show that lines 12-20 run in $\mathcal{O}(k^{4}\log n+k^{2}\nu\log\nu)$ time. Line 12 runs in $\mathcal{O}(k^{2}\nu)$ time according to ˜7.2. Line 13 (calling Algorithm˜1) runs in $\mathcal{O}(k^{2}\nu\log\nu)$ time according to Lemma˜2.5. Line 16 (calling Algorithm˜2) runs in $\mathcal{O}(k^{4}\log n+k\nu)$ time according to Lemma˜6.3. The other lines can be implemented in $\mathcal{O}(k)$ time using hashing. Now, we bound the total run time of lines 7-20 as
@@
Line 22 (calling Algorithm˜3) runs in $\mathcal{O}(k^{2}m\log^{2}n)$ time according to Lemma˜6.9 and does not change the overall time complexity. ∎
Proof of Lemma˜7.3.
Let $S$ be an unbalanced $k$-shredder. We have that $S$ is captured by Algorithm˜4 with probability $1-n^{-99}$ by Lemma˜7.4. By Lemma˜7.5, if $S$ is captured and low-degree, it is put in $\mathcal{L}$ with probability $1$. Otherwise, if $S$ is high-degree, it is put in $\mathcal{L}$ with probability $1-n^{-100}$. Therefore, $S\in\mathcal{L}$ with probability at least $(1-n^{-99})\cdot(1-n^{-100})\geq 1-n^{-98}$ for all $n\geq 2$. Furthermore, every set in $\mathcal{L}$ is a $k$-shredder by Lemma˜7.6. Finally, Algorithm˜4 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ time by Lemma˜7.7. ∎
@@ 8 Listing All $k$-Shredders
Any $k$-shredder can be classified as either balanced or unbalanced. We have not explicitly presented an algorithm for listing balanced $k$-shredders, but it can be found in the appendix (see Algorithm˜7). Combining the two algorithms, we can lists balanced and unbalanced $k$-shredders by calling them sequentially.
@@
1: $G=(V,E)$ - an undirected $k$-vertex-connected graph
2: $\mathcal{L}$ - a list of all $k$-shredders of $G$
3: $\mathcal{L}\leftarrow\varnothing$ 4:
5: $\mathcal{L}_{balanced}\leftarrow$ call Algorithm˜7 on input $G$
6:
$\mathcal{L}\leftarrow\mathcal{L}\cup\mathcal{L}_{balanced}$
7:
8: $\mathcal{L}_{unbalanced}\leftarrow$ call Algorithm˜4 on input $G$
9:
$\mathcal{L}\leftarrow\mathcal{L}\cup\mathcal{L}_{unbalanced}$
10:
11: return $\mathcal{L}$
@@
Let $S$ be a $k$-shredder. Then, $S$ is returned by Algorithm˜5 with probability $1-n^{-98}$. Furthermore, Algorithm˜5 only returns $k$-shredders.
@@
We use proof by casework. If $S$ is balanced, then Lemma˜2.3 implies $S$ is listed by Algorithm˜7 with probability $1-n^{-100}$ (line 6). Otherwise, Lemma˜7.3 implies $S$ is listed by Algorithm˜4 with probability $1-n^{-98}$ (line 9). Hence, with probability $1-n^{-98}$, $S$ is returned by Algorithm˜5. ∎
@@
All $k$-shredders are listed by Algorithm˜5 with probability $1-n^{-97}$.
@@
Jordán proved that there are at most $n$ $k$-shredders in [Jor95]. Lemma˜8.2 implies that each $k$-shredder is listed with probability $1-n^{-98}$. This means that the probability of any one $k$-shredder fails to be listed is at most $n^{-98}$. Using the union bound, we have that the probability that any $k$-shredder fails to be listed is at most $n\cdot n^{-98}=n^{-97}$. Therefore, with probability $1-n^{-97}$, all $k$-shredders are listed. ∎
@@
Algorithm˜5 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ time.
@@
Line 6 runs in $\mathcal{O}(k^{2}m\log n)$ time by Lemma˜2.3. Line 9 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ by Lemma˜7.3. The running time of the last step dominates. ∎
Proof of Lemma˜8.1.
The correctness of Algorithm˜5 is given by Lemmas˜8.2 and 8.3. The time complexity is given by Lemma˜8.4. ∎
We have now shown and proved the existence of a randomized algorithm that improves upon the runtime of $\texttt{All-$k$-shredders}(\cdot)$. The final step is to include the preprocessing step of sparsifying the graph.
@@
As stated earlier, the algorithm developed in [NI92] can be used to preprocess $G$ such that all $k$-shredders are preserved and $m$ is reduced to at most $\mathcal{O}(kn)$ edges. This follows from [CT99, Proposition 3.3]. We can call the sparsifying algorithm and then Algorithm˜5. The correctness of this procedure directly follows from [CT99, Proposition 3.3] and Lemma˜8.1. For time complexity, the sparsifying algorithm takes $\mathcal{O}(m)$ time. Afterwards, we can use Algorithm˜5 to list all $k$-shredders. This takes $\mathcal{O}(k^{4}m\log^{4}n)=\mathcal{O}(k^{4}(kn)\log^{4}n)=\mathcal{O}(k^{5}n\log^{4}n)$ time. In total, both steps take $\mathcal{O}(m+k^{5}n\log^{4}n)$ time. ∎

### v2

@@ 6.2 High-Degree
Proof of Lemma 6.9.
Correctness and time complexity of Algorithm 3 directly follow from Lemmas 6.10 and 6.11. ∎
@@ 7 Capturing and Listing Unbalanced $k$-Shredders
In the previous sections, we showed an algorithm that takes as input a tuple $(x,\nu,\Pi)$, and lists all $k$-shredders that are captured by $(x,\nu,\Pi)$ as well as an unverified set. We then showed how to resolve unverified sets using casework on the structural properties of $k$-shredders. There is one piece of the puzzle left: the method for capturing unbalanced $k$-shredders. For this, we will leverage random sampling using Lemma 6.7 as well as recent developments in local flow algorithms as in [FNS+20].
@@ 7.1 Leveraging Local Flow Algorithms
Theorem 4.1 of [FNS+20] only states that their algorithm returns a vertex cut. But they also construct the set of paths $\Pi$. Their algorithm is simple, and we will briefly explain it here. First, we perform the standard vertex-splitting reduction and reduce the problem to finding directed edge-disjoint paths instead. To find the first path, we use DFS starting from a vertex $x$ and explore $k\nu$ volume. Then, we sample a random endpoint $y_{1}$ among all explored edges. Note that $y_{1}$ is in $L$ with probability at least $1-1/k$ since $\mathrm{vol}(R)\leq\nu$. We reverse the direction of edges on the path from $x$ to $y_{1}$ in the DFS tree and obtain a “residual” graph. Then, we repeat the process in the residual graph to construct the next path from $x$ to $y_{2}$. After $k$ iterations, the endpoints $y_{1},\dots,y_{k}$ of these $k$ paths are in $L$ with probability $(1-1/k)^{k}\geq 1/4$. The paths in the residual graphs can be decomposed via flow decomposition into $k$ directed edge-disjoint paths in the original graph whose total length is $k^{2}\nu$. Finally, these paths correspond to $k$ openly-disjoint paths by the standard reduction in the beginning.
The output of the algorithm described in 7.2 directly corresponds to Definition 2.4. Let $R$ denote the union of all components in $\mathcal{R}$. Notice that $(C,S,R)$ forms a vertex cut such that $\mathrm{vol}(R)<\frac{m}{k}$. The idea is to obtain a seed vertex $x\in R$ using a linear amount of random samples. For each sample, we can directly apply $\texttt{LocalVC}(\cdot,\cdot)$ to obtain the desired set of paths. Furthermore, we can boost the success rate of the algorithm by repeating it a polylogarithmic number of times. In the following section, we formalize this idea.
@@ 7.2 The Algorithm for Unbalanced $k$-Shredders
We first give a high level outline for listing unbalanced $k$-shredders. Let $S$ be an unbalanced $k$-shredder with partition $(C,\mathcal{R})$. With geometric sampling and Lemma 6.7, we will sample a vertex $x\in\mathcal{R}$ with a volume parameter $\nu$ that satisfies $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. Then, we use 7.2 to obtain a set of $k$ openly-disjoint paths $\Pi$, each starting from $x$ and ending at a vertex in $C$ such that the sum of lengths over all paths is at most $k^{2}\nu$. We now have a tuple $(x,\nu,\Pi)$ that captures $S$. After capturing $S$, we call Algorithm 1 to list $S$ as a $k$-shredder or an unverified set. In the latter case, we can verify whether $S$ is a low-degree $k$-shredder using Algorithm 2. If this verification step fails, $S$ must be a high-degree $k$-shredder. In this case, we can add $S$ to a global list of unverified sets. This list will be processed after all unbalanced $k$-shredders have been captured. Lastly, we can extract all high-degree $k$-shredders from the list of unverified sets using Algorithm 3. Pseudocode describing this process is given in Algorithm 4.
@@
$G=(V,E)$- an undirected $k$-vertex-connected graph $\mathcal{L}$- a list containing all unbalanced $k$-shredders of $G$
$\mathcal{L}\leftarrow\varnothing$
2:$\mathcal{U}\leftarrow\varnothing$
3:$\Phi\leftarrow$ initialize a pairwise connectivity oracle as in [Kos23]
[…]
7:$\Pi\leftarrow\texttt{LocalVC}(x,\nu)$ as in [FNS+20]
8:$(\mathcal{L}_{local},U)\leftarrow$ call Algorithm 1 on input $(x,\nu,\Pi)$
[…]
10:call Algorithm 2 on input $(x,\nu,\Pi,U)$IfAlgorithm 2 returned true
11:$\mathcal{L}\leftarrow\mathcal{L}\cup\{U\}$Else
[…]
13:
14:$\mathcal{L}_{high-degree}\leftarrow$ call Algorithm 3 on input $\mathcal{U}$
15:$\mathcal{L}\leftarrow\mathcal{L}\cup\mathcal{L}_{high-degree}$
16:return $\mathcal{L}$
[…]
@@
Let $S$ be an unbalanced $k$-shredder. Then $S$ is captured by a tuple $(x,\nu,\Pi)$ after line 7 during some iteration of the loop on line 4 with probability $1-n^{-99}$.
@@
Suppose $S$ has partition $(C,\mathcal{R})$. Consider the iteration of the loop on line 4 where $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. By Lemma 6.7, we will sample an edge $(x,y)$ such that $x\in\mathcal{R}$ with probability $1-n^{-100}$ during the for loop on line 5. The only missing piece of the tuple is the set of paths. This piece is effectively solved by 7.2. Let $R$ be the union of all components in $\mathcal{R}$. Notice that $(C,S,R)$ is a vertex cut such that $\mathrm{vol}(R)<\frac{m}{k}$. We can call $\texttt{LocalVC}(x,\nu)$ to obtain a set $\Pi$ of $k$ openly-disjoint paths that start from $x$ and end in $C$ such that the sum of lengths over all paths is at most $k^{2}\nu$ with probability $1/4$. This probability is boosted by the for loop on line 6. So the probability that we fail to obtain such a set of paths over all trials is reduced to $(3/4)^{300\log n}\leq(1/2)^{100\log n}=n^{-100}$. Hence, the joint probability that we sample a vertex $x\in\mathcal{R}$ and obtain the desired set of paths $\Pi$ is at least $(1-n^{-100})^{2}\geq 1-2n^{-100}\geq 1-n^{-99}$ for all $n\geq 2$. ∎
@@
Let $S$ be a $k$-shredder captured by the tuple $(x,\nu,\Pi)$ after line 7. If $S$ is low-degree, then $S\in\mathcal{L}$ after line 11. Otherwise, if $S$ is high-degree, then $S\in\mathcal{L}$ after line 15 with probability $1-n^{-100}$.
@@
Suppose $S$ has partition $(C,\mathcal{R})$ and is captured by $(x,\nu,\Pi)$. If Algorithm 1 returns $S$ as a $k$-shredder, we are immediately done by lines 8-9. So let us assume that $S$ is reported as an unverified set. We have that $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. If $S$ is low-degree, we have that there exists a vertex $s\in S$ such that $\deg(s)\leq\nu$. By Lemma 6.3, Algorithm 2 will return true on input $(x,\nu,\Pi,S)$. Then, $S$ is added to $\mathcal{L}$ by line 11. Otherwise, suppose that $S$ is high-degree. Then, $(x,\nu,\Pi,S)$ is added to the set $\mathcal{U}$ by line 12. By Lemma 6.9, Algorithm 3 will return a list of high-degree $k$-shredders containing $S$ on line 14 with probability $1-n^{-100}$. Hence, $S\in\mathcal{L}$ by line 15 with probability $1-n^{-100}$. ∎
@@
There are only three lines of the algorithm where we add sets to $\mathcal{L}$: 8, 11, 15. Line 8 only adds $k$-shredders due to Lemma 2.5. Line 11 only adds $k$-shredders due to Lemma 6.3. Lastly, line 15 only adds $k$-shredders due to Lemma 6.9. ∎
@@
Algorithm 4 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ time.
@@
Firstly, line 3 runs in $\mathcal{O}(km\log n)$ time by 6.2. We show that lines 7-12 run in $\mathcal{O}(k^{4}\log n+k^{2}\nu\log\nu)$ time. Line 7 runs in $\mathcal{O}(k^{2}\nu)$ time according to 7.2. Line 8 (calling Algorithm 1) runs in $\mathcal{O}(k^{2}\nu\log\nu)$ time according to Lemma 2.5. Line 10 (calling Algorithm 2) runs in $\mathcal{O}(k^{4}\log n+k\nu)$ time according to Lemma 6.3. The other lines can be implemented in $\mathcal{O}(k)$ time using hashing. Now, we bound the total run time of lines 4-12 as
@@
Line 14 (calling Algorithm 3) runs in $\mathcal{O}(k^{2}m\log^{2}n)$ time according to Lemma 6.9 and does not change the overall time complexity. ∎
Proof of Lemma 7.3.
Let $S$ be an unbalanced $k$-shredder. We have that $S$ is captured by Algorithm 4 with probability $1-n^{-99}$ by Lemma 7.4. By Lemma 7.5, if $S$ is captured and low-degree, it is put in $\mathcal{L}$ with probability $1$. Otherwise, if $S$ is high-degree, it is put in $\mathcal{L}$ with probability $1-n^{-100}$. Therefore, $S\in\mathcal{L}$ with probability at least $(1-n^{-99})\cdot(1-n^{-100})\geq 1-n^{-98}$ for all $n\geq 2$. Furthermore, every set in $\mathcal{L}$ is a $k$-shredder by Lemma 7.6. Finally, Algorithm 4 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ time by Lemma 7.7. ∎
@@ 8 Listing All $k$-Shredders
Any $k$-shredder can be classified as either balanced or unbalanced. We have not explicitly presented an algorithm for listing balanced $k$-shredders, but it can be found in the appendix (see Algorithm 7). Combining the two algorithms, we can lists balanced and unbalanced $k$-shredders by calling them sequentially.
@@
$G=(V,E)$- an undirected $k$-vertex-connected graph $\mathcal{L}$- a list of all $k$-shredders of $G$
$\mathcal{L}\leftarrow\varnothing$
2:
3:$\mathcal{L}_{balanced}\leftarrow$ call Algorithm 7 on input $G$
4:$\mathcal{L}\leftarrow\mathcal{L}\cup\mathcal{L}_{balanced}$
5:
6:$\mathcal{L}_{unbalanced}\leftarrow$ call Algorithm 4 on input $G$
7:$\mathcal{L}\leftarrow\mathcal{L}\cup\mathcal{L}_{unbalanced}$
8:
9:return $\mathcal{L}$
[…]
@@
Let $S$ be a $k$-shredder. Then, $S$ is returned by Algorithm 5 with probability $1-n^{-98}$. Furthermore, Algorithm 5 only returns $k$-shredders.
@@
We use proof by casework. If $S$ is balanced, then Lemma 2.3 implies $S$ is listed by Algorithm 7 with probability $1-n^{-100}$ (line 4). Otherwise, Lemma 7.3 implies $S$ is listed by Algorithm 4 with probability $1-n^{-98}$ (line 7). Hence, with probability $1-n^{-98}$, $S$ is returned by Algorithm 5. ∎
@@
All $k$-shredders are listed by Algorithm 5 with probability $1-n^{-97}$.
@@
Jordán proved that there are at most $n$ $k$-shredders in [Jor95]. Lemma 8.2 implies that each $k$-shredder is listed with probability $1-n^{-98}$. This means that the probability of any one $k$-shredder fails to be listed is at most $n^{-98}$. Using the union bound, we have that the probability that any $k$-shredder fails to be listed is at most $n\cdot n^{-98}=n^{-97}$. Therefore, with probability $1-n^{-97}$, all $k$-shredders are listed. ∎
@@
Algorithm 5 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ time.
@@
Line 4 runs in $\mathcal{O}(k^{2}m\log n)$ time by Lemma 2.3. Line 7 runs in $\mathcal{O}(k^{4}m\log^{4}n)$ by Lemma 7.3. The running time of the last step dominates. ∎
Proof of Lemma 8.1.
The correctness of Algorithm 5 is given by Lemmas 8.2 and 8.3. The time complexity is given by Lemma 8.4. ∎
[…]
@@
As stated earlier, the algorithm developed in [NI92] can be used to preprocess $G$ such that all $k$-shredders are preserved and $m$ is reduced to at most $\mathcal{O}(kn)$ edges. This follows from [CT99, Proposition 3.3]. We can call the sparsifying algorithm and then Algorithm 5. The correctness of this procedure directly follows from [CT99, Proposition 3.3] and Lemma 8.1. For time complexity, the sparsifying algorithm takes $\mathcal{O}(m)$ time. Afterwards, we can use Algorithm 5 to list all $k$-shredders. This takes $\mathcal{O}(k^{4}m\log^{4}n)=\mathcal{O}(k^{4}(kn)\log^{4}n)=\mathcal{O}(k^{5}n\log^{4}n)$ time. In total, both steps take $\mathcal{O}(m+k^{5}n\log^{4}n)$ time. ∎